You tie a 5.8 kg mass to the end of a 88 cm long string. You then swing the mass a horizontal circle above your head until the string breaks. If you know the string can withstand a maximum tension of 1,503 N without breaking, what was the linear speed (in m/s) of the mass when the string broke?
1
Expert's answer
2020-06-12T11:15:51-0400

The string breaks when the tension becomes higher than the allowable. The tension is caused by the centrifugal force:


"T=\frac{mv^2}{R},\\\space\\
v=\sqrt{\frac{TR}{m}}=\sqrt{\frac{1.503\cdot0.88}{5.8}}=0.48\text{ m/s}."